# LongLLaDA: Unlocking Long Context Capabilities in Diffusion LLMs

### v2

@@ 1 Introduction
Recently, diffusion LLMs have become widely discussed in Natural Language Processing research (Nie et al., 2025; Ye et al., 2025). They are regarded as a potential solution to key limitations of traditional auto-regressive LLMs (Touvron et al., 2023; Sun et al., 2024), including the reversal curse (Berglund et al., 2023), complex reasoning (Dziri et al., 2023), and maintaining coherence across extended contexts (Bachmann & Nagarajan, 2024; Ye et al., 2024; 2025). Significant research efforts have focused on validating their scalability (Nie et al., 2025; Ye et al., 2025), adapting them for multimodality (Yang et al., 2025; You et al., 2025; Yu et al., 2025), applying them to reasoning tasks (Zhao et al., 2025; Huang et al., 2025; Zhu et al., 2025), and optimizing their efficiency (Ma et al., 2025; Hu et al., 2025; Wu et al., 2025). However, the long-context capabilities of diffusion LLMs, specifically their performance and potential for length extrapolation, remain unexplored.
We begin by systematically evaluating diffusion LLM LLaDA (Nie et al., 2025) against auto-regressive LLM LLaMA3 (Meta, 2024a) on perplexity and retrieval tasks, both within and beyond their pretrained context lengths (Figure 1). Notably, diffusion LLMs maintain stable perplexity and exhibit localized perception during direct length extrapolation. In stark contrast, auto-regressive LLMs suffer catastrophic perplexity surges and performance collapse when input length exceeds their maximum supported context window, 8k tokens. This divergence reveals fundamental architectural differences in long-context handling, raising critical questions: (1) What mechanisms enable diffusion LLMs’ extrapolation stability? (2) Can established length-extension techniques for auto-regressive LLMs be transferred to diffusion architectures? (3) How do diffusion LLMs perform on long-context benchmarks relative to auto-regressive baselines, and what unique capabilities or limitations emerge?
In this work, we address these questions through comprehensive experiments and analysis. Besides the perplexity and retrieval experiment, we also benchmark Needle-In-A-Haystack (NIAH) performance for diffusion LLMs (LLaDA (Nie et al., 2025), LLaDA-1.5 (Zhu et al., 2025), Dream-v0 (Ye et al., 2025)), quantitatively confirming their local perception bias during length extrapolation. We then analyze this phenomenon through Rotary Position Embedding (RoPE) theory, validating our interpretation with t-SNE visualizations. Building on these insights, we propose LongLLaDA, a training-free method which successfully extends LLaDA’s context window using NTK-based RoPE extrapolation (bloc97, 2023b), and verify preserved scaling laws (Liu et al., 2023b). Finally, we identify task-dependent capabilities where diffusion LLMs surpass or lag behind auto-regressive counterparts on long-context benchmarks. Our contributions are summarized as follows:
  • Effective context extension demonstrating NTK-based RoPE extrapolation and scaling laws transfer seamlessly to diffusion LLMs, achieving 6$\times$ context expansion (24k tokens) without further training.
  • Capability benchmarking revealing diffusion LLMs match auto-regressive models on retrieval tasks, lag in aggregation, but excel at QA. We provide foundational insights for future long-context diffusion research.
@@ 2 Long-Context Phenomenology of Diffusion LLMs
Figure 2: Results of Needle-In-A-Haystack tests (Gkamradt, 2023) on LLaDA-8B Series (Nie et al., 2025) and LLaMA3-8B Series (Meta, 2024b) under direct extrapolation.
@@
Figure 3: NIAH Results of LLaDA-8B-Base (Nie et al., 2025) with different sampling steps $s$.
We first evaluate the length extrapolation capabilities of diffusion LLMs, including LLaDA (Nie et al., 2025), LLaDA-1.5 (Zhu et al., 2025), and Dream-v0 (Ye et al., 2025), compared with auto-regressive LLMs such as LLaMA3 (Meta, 2024a), via Needle-In-A-Haystack (Gkamradt, 2023; Li et al., 2024), based on the experimental setup in Appendix B.1. All LLMs are required to generate at most 32 tokens, with diffusion LLMs using a block size and sampling steps of 32. The results are shown in Figure 2. LLaMA3-8B-Base and LLaMA3-8B-Instruct maintain perfect retrieval accuracy within their pretrained 8k length, but suffer catastrophic performance degradation beyond this limit, failing to retrieve information at any depth. In contrast, LLaDA-8B-Base and LLaDA-8B-Instruct achieve 100% retrieval accuracy within a 4k context. Surprisingly, when exceeding 4k, up to 24k, LLaDA still retrieves information from the nearest 4k window, demonstrating a local perception like a sliding window. This behavior remarkably differs from auto-regressive LLM extrapolation. Similar phenomena are observed in LLaDA-1.5 and Dream-v0, as illustrated in Appendix B.2.
@@ 3 Mechanistic Analysis
According to the preliminary knowledge in Appendix A, we attribute this phenomenon to diffusion LLMs being trained with richer positional information compared to auto-regressive LLMs. Critically, the bidirectional attention mechanisms in diffusion LLMs expose them to relative position rage of $[1-T_{\text{train}},T_{\text{train}}-1]$ during training, contrasting with the $[0,T_{\text{train}}-1]$ range typical of auto-regressive models. This difference is evident in the RoPE mechanism. As visualized in Figure 4, for LLaDA ($T_{\text{train}}=$ 4k) and LLaMA ($T_{\text{train}}=$ 8k), we observe how the positional embeddings (sine/cosine components) behave within and beyond their maximum trained relative positions.
@@
  • Moderate Frequencies: LLaMA3’s auto-regressive attention observes relative positions $[0,8191]$ when trained on 8192-token sequences. In contrast, LLaDA’s bidirectional attention observes symmetric relative positions $[-4095,4095]$ despite its shorter 4096-token training length. This symmetric coverage provides a key advantage by fully capturing a complete period of both the cosine and sine, enhancing its tolerance of direct length extrapolation.
@@
This results in a relatively flattened perplexity growth curve, similar to auto-regressive RoPE-based LLMs with a smaller base (Liu et al., 2023b; Men et al., 2024), as detailed in Appendix A. However, since the cosine function in RoPE, which primarily captures relative distances, is even, negative relative positions do not increase the LLM’s maximum perceivable distance in the pre-training stage. Thus, diffusion LLM can only retrieve key information from limited relative positions within the training length, leading to the observed decay pattern in the NIAH evaluation.
We validate this interpretation with the t-SNE visualization (Van der Maaten & Hinton, 2008; Zandieh et al., 2024) of QK states from the final layer of LLaMA3-8B-Base (Meta, 2024a) and LLaDA-8B-Base (Nie et al., 2025), as shown in Figure 5. As shown in Figure 5(a), for auto-regressive LLMs such as LLaMA3-8B-Base, the QK states within and beyond the maximum supported context length, 8k, present two different distribution clusters, and the manifold for QK states with RoPE also shows a different trend when position embedding becomes OOD. Comparatively, regarding the clustering feature for diffusion LLMs such as LLaDA-8B-Base, there is no distribution shift between QK states within and beyond 4k, and a uniform manifold for QK states with RoPE. This demonstrates that diffusion LLM is more robust for the OOD position embeddings in length extrapolation. Therefore, unlike traditional auto-regressive LLMs that exhibit catastrophic performance degradation when exceeding their maximum supported context length, diffusion LLMs maintain stable outputs and demonstrate local perception in extended context.
@@
Figure 5: Visualization of the QK states from the final layer of LLaMA3-8B-Base (Meta, 2024a) and LLaDA-8B-Base (Nie et al., 2025) for sample from the GovReport subsets in LongBench (Bai et al., 2023). The visualization uses a 2D t-SNE projection (Van der Maaten & Hinton, 2008), with each token represented as a point in the image and the position index shown via color changing.
@@ 4 Context Extension For Diffusion LLMs
Since the reason for the surprising phenomenon has been clarified, we now move on to the extrapolation methods for diffusion LLMs. Since the retrievable depth of diffusion LLMs remains constrained by the range of cosine values encountered during pre-training, we transfer the NTK-based extrapolation (bloc97, 2023b) and its scaling laws (Liu et al., 2023b) to diffusion LLMs, thus proposing the length extrapolation method for diffusion LLMs, LongLLaDA. As detailed in Appendix A. The scaling factor $\lambda$ in training-free NTK scaling (bloc97, 2023b) for RoPE-based auto-regressive LLMs is decided by the extrapolation context length $t$ and critical dimension $d_{\text{extra}}$ calculated by rotary base $\beta_{0}$ and pretrained context length $T_{\text{train}}$, as shown in Equation 1.
@@
Similarly, in LongLLaDA, based on Nie et al. (2025), the pretrained rotary base $\beta_{0}=500000$, and the pre-training context length $T_{\text{train}}$ is 4k. This yields a critical dimension $d_{\text{extra}}=64$. Accordingly, the required scaling factor $\lambda$ for extrapolation to 8k, 16k, 24k, and 32k is calculated as 4, 14, 31, and 55, respectively. The extrapolation results are illustrated in Figure 6 and Figure 7.
When $\lambda=4,14$, LongLLaDA can effectively extrapolate diffusion LLMs to the corresponding context lengths, achieving near 100% recall across all depths within these ranges. As the context length increases beyond the extrapolation limit, the retrievable depth proportionally expands while maintaining the local-perception effect. The average depth score curves exhibit a right shift across different context lengths. When $\lambda=31$, a lost-in-the-middle phenomenon (Liu et al., 2023a) similar to auto-regressive models emerges in intermediate depths, indicating that LongLLaDA approaches its practical extrapolation limit (bloc97, 2023b). When $\lambda=55$, further extrapolation is unachievable. We also validate the effectiveness of LongLLaDA in LLaDA-1.5 (Zhu et al., 2025) and Dream-v0 (Ye et al., 2025) in Appendix B.2. Consequently, for RoPE-based diffusion LLMs, NTK extrapolation and its scaling law remain applicable during inference.
@@
Figure 6: NIAH Results of LLaDA-8B-Base (Nie et al., 2025) with different RoPE scaling factor.
@@
Figure 7: NIAH Results of LLaDA-8B-Instruct (Nie et al., 2025) with different RoPE scaling factor.
@@ 5 Task-Driven Long-Context Capability Analysis
Table 1: Results of LLaDA-8B (Nie et al., 2025), LLaDA-1.5 (Zhu et al., 2025) and LLaMA3-8B (Meta, 2024b) on LongBench (Bai et al., 2023) under 4k and 8k context length. Gray cells indicate that the evaluation context length exceeds the context length supported by the evaluated LLM. SD, MD, Sum, and Syn stand for Single-Doc QA, Multi-Doc QA, Summarization, and Synthetic tasks, while Avg is the average score of all subtasks weighted by the evaluation data number.
@@
+ NTK $\lambda=4$ 99.5 82.3 80.5 92.6 96.4 61.0 73.0 84.7 51.8 17.1 53.5 44.1
+ NTK $\lambda=14$ 99.8 83.5 77.0 92.5 99.3 68.9 64.0 86.8 85.4 48.1 54.0 72.0
+ NTK $\lambda=31$ 100.0 83.8 77.0 92.7 97.8 75.2 62.5 87.1 97.2 51.8 41.0 78.0
@@
+ NTK $\lambda=4$ 99.8 65.5 89.5 90.3 95.9 56.6 89.0 85.8 41.6 31.0 73.0 44.0
+ NTK $\lambda=14$ 100.0 76.4 89.0 92.9 97.3 66.2 89.5 88.9 67.1 53.8 84.5 66.7
+ NTK $\lambda=31$ 100.0 77.7 86.5 92.8 98.8 73.5 88.5 91.3 88.0 62.2 82.0 81.1
@@
+ NTK $\lambda=4$ 99.8 73.9 91.0 92.5 96.3 59.3 88.0 86.5 43.3 32.2 73.0 45.3
+ NTK $\lambda=14$ 100.0 79.8 88.5 93.6 99.9 67.8 89.0 90.8 67.4 51.6 84.0 66.3
+ NTK $\lambda=31$ 100.0 81.6 87.5 93.8 98.9 75.1 86.5 91.5 85.8 58.2 81.5 78.7
@@
Table 2: Results of LLaDA-8B (Nie et al., 2025), LLaDA-1.5 (Zhu et al., 2025) and LLaMA3-8B (Meta, 2024b) on RULER (Hsieh et al., 2024) under 4k, 8k and 16k context length.
@@
Regarding the downstream long-context performance of diffusion LLMs and their difference from traditional auto-regressive LLMs, apart from the NIAH retrieval evaluation, we conduct comparative analyses across more benchmarks using LLaDA and LLaMA as examples. We first evaluate LLaDA-8B (Nie et al., 2025), LLaDA-1.5 (Zhu et al., 2025), and LLaMA3-8B (Meta, 2024a), including pre-trained models and those employing NTK-based extrapolation during inference, with LongBench (Bai et al., 2023), in 4k and 8k context length, with the exceeding part being truncated from the middle. For the summary tasks, the output length is 512, while for the others, the output length is 64. We still keep the sampling steps the same as the output length, and the block size to 64 for diffusion LLMs. The results are shown in Table 1. Still, LLaDA can give a stable output and get a decent performance beyond the maximum supported context length. Moreover, we find that in all task domains besides synthetic tasks, the difference between LLaDA Series and LLaMA3 Series is relatively limited compared with the difference within LLaMA3 Series. Only in the synthetic domain do LLaDA Series outperform LLaMA3 Series consistently. This inspires us to conduct an in-depth discussion of diffusion LLMs on the performance of the synthesis tasks compared with auto-regressive LLMs.
@@
We further the discussion with RULER benchmark (Hsieh et al., 2024), we compare LLaDA-8B (Nie et al., 2025), LLaDA-1.5 (Zhu et al., 2025), and LLaMA3-8B (Meta, 2024a), at context lengths of 4k, 8k, and 16k. We set the block size and sampling steps to 64 for diffusion LLMs. The results are shown in Table 2. First, consistent with the NIAH results, auto-regressive LLMs fail to produce valid outputs beyond their effective context length, while diffusion LLMs maintain measurable performance. Regarding task types, diffusion LLMs achieve comparable results to auto-regressive LLMs on NIAH tasks, including Single-Key, Multi-Key, Multi-Query, and Multi-Value variants. However, diffusion LLMs show significantly inferior performance in aggregation tasks, including Variable Tracing and Frequent or Common Word Extraction, where auto-regressive LLMs typically perform well. Surprisingly, on QA tasks, including SQuAD and Hotpot, that challenge auto-regressive LLMs (Hsieh et al., 2024), diffusion LLMs demonstrate superior capability. These observations reveal the distinctive characteristics of diffusion LLMs in long-context tasks, that current diffusion LLMs, like LLaDA, demonstrate comparable performance to the auto-regressive LLMs, like LLaMA3, in most task types, but underperform in aggregation tasks, and outperform in QA tasks consistently.

### v3

@@ 1 Introduction
Recently, diffusion LLMs have become widely discussed in Natural Language Processing research (nie2025large; dream2025). They are regarded as a potential solution to key limitations of traditional auto-regressive LLMs (touvron2023llama; Sun2024MOSS), including the reversal curse (berglund2023reversal), complex reasoning (dziri2023faith), and maintaining coherence across extended contexts (bachmann2024pitfalls; ye2024beyond; dream2025). Significant research efforts have focused on validating their scalability (nie2025large; dream2025), adapting them for multimodality (yang2025mmada; you2025llada; yu2025dimple), applying them to reasoning tasks (zhao2025d1; huang2025reinforcing; zhu2025llada), and optimizing their efficiency (ma2025dkv; hu2025accelerating; wu2025fast). However, the long-context capabilities of diffusion LLMs, specifically their performance and potential for length extrapolation, remain unexplored.
We begin by systematically evaluating diffusion LLM LLaDA (nie2025large) against auto-regressive LLM LLaMA3 (meta2024introducing) on perplexity and retrieval tasks, both within and beyond their pretrained context lengths (Figure 1). Notably, diffusion LLMs maintain stable perplexity and exhibit localized perception during direct length extrapolation. In stark contrast, auto-regressive LLMs suffer catastrophic perplexity surges and performance collapse when input length exceeds their maximum supported context window, 8k tokens. This divergence reveals fundamental architectural differences in long-context handling, raising critical questions: (1) What mechanisms enable diffusion LLMs’ extrapolation stability? (2) Can established length-extension techniques for auto-regressive LLMs be transferred to diffusion architectures? (3) How do diffusion LLMs perform on long-context benchmarks relative to auto-regressive baselines, and what unique capabilities or limitations emerge?
In this work, we address these questions through comprehensive experiments and analysis. Besides the perplexity and retrieval experiment, we also benchmark Needle-In-A-Haystack (NIAH) performance for diffusion LLMs (LLaDA (nie2025large), LLaDA-1.5 (zhu2025llada), Dream-v0 (dream2025)), quantitatively confirming their local perception bias during length extrapolation. We then analyze this phenomenon through Rotary Position Embedding (RoPE) theory, validating our interpretation with t-SNE visualizations. Building on these insights, we propose LongLLaDA, a training-free method which successfully extends LLaDA’s context window using NTK-based RoPE extrapolation (fixedNTK), and verify preserved scaling laws (liu2023scaling). Finally, we identify task-dependent capabilities where diffusion LLMs surpass or lag behind auto-regressive counterparts on long-context benchmarks. Our contributions are summarized as follows:
[…]
  • Capability benchmarking revealing diffusion LLMs match auto-regressive models on retrieval tasks, lag in aggregation, but excel at synthetic QA. We provide foundational insights for future long-context diffusion research.
@@ 2 Long-Context Phenomenology of Diffusion LLMs
Figure 2: Results of Needle-In-A-Haystack tests (needle_in_a_haystack) on LLaDA-8B Series (nie2025large) and LLaMA3-8B Series (meta2024llama) under direct extrapolation.
@@
Figure 3: NIAH Results of LLaDA-8B-Base (nie2025large) with different sampling steps $s$.
We first evaluate the length extrapolation capabilities of diffusion LLMs, including LLaDA (nie2025large), LLaDA-1.5 (zhu2025llada), and Dream-v0 (dream2025), compared with auto-regressive LLMs such as LLaMA3 (meta2024introducing), via Needle-In-A-Haystack (needle_in_a_haystack; li2024needlebench), based on the experimental setup in Appendix B.1. All LLMs are required to generate at most 32 tokens, with diffusion LLMs using a block size and sampling steps of 32. The results are shown in Figure 2. LLaMA3-8B-Base and LLaMA3-8B-Instruct maintain perfect retrieval accuracy within their pretrained 8k length, but suffer catastrophic performance degradation beyond this limit, failing to retrieve information at any depth. In contrast, LLaDA-8B-Base and LLaDA-8B-Instruct achieve 100% retrieval accuracy within a 4k context. Surprisingly, when exceeding 4k, up to 24k, LLaDA still retrieves information from the nearest 4k window, demonstrating a local perception like a sliding window. This behavior remarkably differs from auto-regressive LLM extrapolation. Similar phenomena are observed in LLaDA-1.5 and Dream-v0, as illustrated in Appendix B.2.
@@ 3 Mechanistic Analysis
According to the preliminary knowledge in Appendix A, we attribute this phenomenon to diffusion LLMs being trained with richer positional information compared to auto-regressive LLMs. Critically, the bidirectional attention mechanisms in diffusion LLMs expose them to relative position rage of $1-T_{\text{train}},T_{\text{train}}-1]$ during training, contrasting with the $0,T_{\text{train}}-1]$ range typical of auto-regressive models. This difference is evident in the RoPE mechanism. As visualized in Figure 4, for LLaDA ($T_{\text{train}}=$ 4k) and LLaMA ($T_{\text{train}}=$ 8k), we observe how the positional embeddings (sine/cosine components) behave within and beyond their maximum trained relative positions.
@@
  • Moderate Frequencies: LLaMA3’s auto-regressive attention observes relative positions $[0,8191]$ when trained on 8192-token sequences. In contrast, LLaDA’s bidirectional attention observes symmetric relative positions [4095,4095][-4095,4095] despite its shorter 4096-token training length. This symmetric coverage provides a key advantage by fully capturing a complete period of both the cosine and sine, enhancing its tolerance of direct length extrapolation.
@@
This results in a relatively flattened perplexity growth curve, similar to auto-regressive RoPE-based LLMs with a smaller base (liu2023scaling; men2024base), as detailed in Appendix A. However, since the cosine function in RoPE, which primarily captures relative distances, is even, negative relative positions do not increase the LLM’s maximum perceivable distance in the pre-training stage. Thus, diffusion LLM can only retrieve key information from limited relative positions within the training length, leading to the observed decay pattern in the NIAH evaluation.
We validate this interpretation with the t-SNE visualization (van2008visualizing; zandieh2024subgen) of QK states from the final layer of LLaMA3-8B-Base (meta2024introducing) and LLaDA-8B-Base (nie2025large), as shown in Figure 5. As shown in Figure 5(a), for auto-regressive LLMs such as LLaMA3-8B-Base, the QK states within and beyond the maximum supported context length, 8k, present two different distribution clusters, and the manifold for QK states with RoPE also shows a different trend when position embedding becomes OOD. Comparatively, regarding the clustering feature for diffusion LLMs such as LLaDA-8B-Base, there is no distribution shift between QK states within and beyond 4k, and a uniform manifold for QK states with RoPE. This demonstrates that diffusion LLM is more robust for the OOD position embeddings in length extrapolation. Therefore, unlike traditional auto-regressive LLMs that exhibit catastrophic performance degradation when exceeding their maximum supported context length, diffusion LLMs maintain stable outputs and demonstrate local perception in extended context.
@@
Figure 5: Visualization of the QK states from the final layer of LLaMA3-8B-Base (meta2024introducing) and LLaDA-8B-Base (nie2025large) for sample from the GovReport subsets in LongBench (bai2023longbench). The visualization uses a 2D t-SNE projection (van2008visualizing), with each token represented as a point in the image and the position index shown via color changing.
@@ 4 Context Extension For Diffusion LLMs
Since the reason for the surprising phenomenon has been clarified, we now move on to the extrapolation methods for diffusion LLMs. Since the retrievable depth of diffusion LLMs remains constrained by the range of cosine values encountered during pre-training, we transfer the NTK-based extrapolation (fixedNTK) and its scaling laws (liu2023scaling) to diffusion LLMs, thus proposing the length extrapolation method for diffusion LLMs, LongLLaDA. As detailed in Appendix A. The scaling factor $\lambda$ in training-free NTK scaling (fixedNTK) for RoPE-based auto-regressive LLMs is decided by the extrapolation context length $t$ and critical dimension $d_{\text{extra}}$ calculated by rotary base $\beta_{0}$ and pretrained context length $T_{\text{train}}$, as shown in Equation 1.
@@
Similarly, in LongLLaDA, based on nie2025large, the pretrained rotary base $\beta_{0}=500000$, and the pre-training context length $T_{\text{train}}$ is 4k. This yields a critical dimension $d_{\text{extra}}=64$. Accordingly, the required scaling factor $\lambda$ for extrapolation to 8k, 16k, 24k, and 32k is calculated as 4, 14, 31, and 55, respectively. The extrapolation results are illustrated in Figure 6 and Figure 7.
When $\lambda=4,14$, LongLLaDA can effectively extrapolate diffusion LLMs to the corresponding context lengths, achieving near 100% recall across all depths within these ranges. As the context length increases beyond the extrapolation limit, the retrievable depth proportionally expands while maintaining the local-perception effect. The average depth score curves exhibit a right shift across different context lengths. When $\lambda=31$, a lost-in-the-middle phenomenon (liu2023lost) similar to auto-regressive models emerges in intermediate depths, indicating that LongLLaDA approaches its practical extrapolation limit (fixedNTK). When $\lambda=55$, further extrapolation is unachievable. We also validate the effectiveness of LongLLaDA in LLaDA-1.5 (zhu2025llada) in Appendix B.2. Consequently, for RoPE-based diffusion LLMs, NTK extrapolation and its scaling law remain applicable during inference.
@@
Figure 6: NIAH Results of LLaDA-8B-Base (nie2025large) with different RoPE scaling factor.
@@
Figure 7: NIAH Results of LLaDA-8B-Instruct (nie2025large) with different RoPE scaling factor.
@@ 5 Task-Driven Long-Context Capability Analysis
Table 1: Results of LLaDA-8B (nie2025large), LLaDA-1.5 (zhu2025llada) and LLaMA3-8B (meta2024llama) on LongBench (bai2023longbench) under 4k and 8k context length. Gray cells indicate that the evaluation context length exceeds the context length supported by the evaluated LLM. SD, MD, Sum, and Syn stand for Single-Doc QA, Multi-Doc QA, Summarization, and Synthetic tasks, while Avg is the average score of all subtasks weighted by the evaluation data number.
@@
+ $\lambda=4$ 99.5 82.3 80.5 92.6 96.4 61.0 73.0 84.7 51.8 17.1 53.5 44.1
+ $\lambda=14$ 99.8 83.5 77.0 92.5 99.3 68.9 64.0 86.8 85.4 48.1 54.0 72.0
+ $\lambda=31$ 100.0 83.8 77.0 92.7 97.8 75.2 62.5 87.1 97.2 51.8 41.0 78.0
@@
+ $\lambda=4$ 99.8 65.5 89.5 90.3 95.9 56.6 89.0 85.8 41.6 31.0 73.0 44.0
+ $\lambda=14$ 100.0 76.4 89.0 92.9 97.3 66.2 89.5 88.9 67.1 53.8 84.5 66.7
+ $\lambda=31$ 100.0 77.7 86.5 92.8 98.8 73.5 88.5 91.3 88.0 62.2 82.0 81.1
@@
+ $\lambda=4$ 99.8 73.9 91.0 92.5 96.3 59.3 88.0 86.5 43.3 32.2 73.0 45.3
+ $\lambda=14$ 100.0 79.8 88.5 93.6 99.9 67.8 89.0 90.8 67.4 51.6 84.0 66.3
+ $\lambda=31$ 100.0 81.6 87.5 93.8 98.9 75.1 86.5 91.5 85.8 58.2 81.5 78.7
@@
Table 2: Results of LLaDA-8B (nie2025large), LLaDA-1.5 (zhu2025llada) and LLaMA3-8B (meta2024llama) on RULER (hsieh2024ruler) under 4k, 8k and 16k context length.
@@
Regarding the downstream long-context performance of diffusion LLMs and their difference from traditional auto-regressive LLMs, apart from the NIAH retrieval evaluation, we conduct comparative analyses across more benchmarks using LLaDA and LLaMA as examples. We first evaluate LLaDA-8B (nie2025large), LLaDA-1.5 (zhu2025llada), and LLaMA3-8B (meta2024introducing), including pre-trained models and those employing NTK-based extrapolation during inference, with LongBench (bai2023longbench), in 4k and 8k context length, with the exceeding part being truncated from the middle. For the summary tasks, the output length is 512, while for the others, the output length is 64. We still keep the sampling steps the same as the output length, and the block size to 64 for diffusion LLMs. The results are shown in Table 1. Still, LLaDA can give a stable output and get a decent performance beyond the maximum supported context length. Moreover, we find that in all task domains besides synthetic tasks, the difference between LLaDA Series and LLaMA3 Series is relatively limited compared with the difference within LLaMA3 Series. Only in the synthetic domain do LLaDA Series outperform LLaMA3 Series consistently. This inspires us to conduct an in-depth discussion of diffusion LLMs on the performance of the synthesis tasks compared with auto-regressive LLMs.
@@
We further the discussion with RULER benchmark (hsieh2024ruler), we compare LLaDA-8B (nie2025large), LLaDA-1.5 (zhu2025llada), and LLaMA3-8B (meta2024introducing), at context lengths of 4k, 8k, and 16k. We set the block size and sampling steps to 64 for diffusion LLMs. The results are shown in Table 2. First, consistent with the NIAH results, auto-regressive LLMs fail to produce valid outputs beyond their effective context length, while diffusion LLMs maintain measurable performance. Regarding task types, diffusion LLMs achieve comparable results to auto-regressive LLMs on NIAH tasks, including Single-Key, Multi-Key, Multi-Query, and Multi-Value variants. However, diffusion LLMs show significantly inferior performance in aggregation tasks, including Variable Tracing and Frequent or Common Word Extraction, where auto-regressive LLMs typically perform well. Surprisingly, on QA tasks, including SQuAD and Hotpot, that challenge auto-regressive LLMs (hsieh2024ruler), diffusion LLMs demonstrate superior capability. These observations reveal the distinctive characteristics of diffusion LLMs in long-context tasks, that current diffusion LLMs, like LLaDA, demonstrate comparable performance to the auto-regressive LLMs, like LLaMA3, in most task types, but underperform in aggregation tasks, and outperform in synthetic QA tasks consistently.